# Decoding Generalization from Memorization in Deep Neural Networks

### v1

@@ A.4 Comparison of decoding results and early stopping test accuracies
A.5 Experiments with AlexNet model trained on Tiny ImageNet
@@
MASC test accuracy over the layers of AlexNet trained on Tiny ImageNet when the data is projected onto corrupted training subspaces with the indicated corruption degree is shown in Figure 7. Testing accuracy of the model and best model testing accuracy is shown for comparison. Best Model Testing Accuracy corresponds accuracy of the testing data of the model if early stopping was used.
@@
Even for AlexNet-Tiny ImageNet corrupted model (with non-zero corruption degree), except those with 100% corruption degree, we find that our Minimum Angle Subspace Classifier (MASC) in at least one layer has better testing accuracy than the corresponding model itself. In Table 4, the MASC accuracy on testing for the best layer as well as by what percentage the MASC classifier outperformed the model for the best layer for each corruption degrees 20%, 40%, 60% and 80% is documented.
We have also performed MASC with 90% variance explained for PCA on AlexNet trained on Tiny ImageNet. The comparison between 99% and 90% variance captured for PCA is shown in Section A.12.
MASC test accuracy over the layers of AlexNet trained on Tiny ImageNet when the data set is projected onto corrupted training and true training subspace is shown in Figure 8. In Table 5, the MASC accuracy on testing for the best layer as well as by what percentage the MASC classifier outperformed the model for the best layer for each corruption degrees 20%, 40%, 60% and 80% is documented.
We have also performed an addition experiments with 99.9% PCA variance explained on AlexNet trained on Tiny ImageNet. We find that with high percentage of PCA variance explained, the MASC performs better on training dataset with true labels. The results for MASC of 99.9% , 99% and 90% variance explained are shown in Figure 38.
Table 4: Subspace constructed using corrupted labels - Percentage the MASC classifier outperformed the model (AlexNet-Tiny ImageNet).
Corruption degree 20% 40% 60% 80%
Model Test Accuracy (%) 9.74 5.44 2.02 0.73
MASC Accuacy on Testing (Best layer) 12.42 8.36 2.93 0.83
MASC outperformed the model (%) 27.51 53.67 45.04 13.69
Figure 7: MASC test accuracy over the layers of AlexNet trained on Tiny ImageNet when the data is projected onto corrupted training subspaces with the indicated corruption degree. Testing accuracy of the model and best model testing accuracy is shown for comparison. Best Model Testing Accuracy corresponds accuracy of the testing data of the model if early stopping was used.
@@
Figure 8: MASC test accuracy over the layers of AlexNet trained on Tiny ImageNet when the data set is projected onto corrupted training and true training subspace. Testing accuracy of the model and best model testing accuracy is shown for comparison. Best Model Testing Accuracy corresponds accuracy of the testing data of the model if early stopping was used.
Table 5: Subspace constructed using true labels - Percentage the MASC classifier outperformed the model (AlexNet-Tiny ImageNet).
Corruption degree 20% 40% 60% 80%
Model Test Accuracy (%) 9.74 5.44 2.02 0.73
MASC Accuacy on Testing (Best layer) 12.99 10.02 5.19 2.27
MASC outperformed the model (%) 33.36 84.19 156.93 210.95
A.6 Randomly initialized control models vs Trained models
@@
This section covers a set of control experiments to show MASC performance on random initialized model and contrast this with the trained models presented in the main text. We have verified that, for every such control model, the model training and testing accuracies for the randomly initialized models is at chance level, for the corresponding dataset in question.
@@
MASC accuracy on testing for randomly initialized model and trained model when data is projected on corrupted training subspaces is shown in Figure 9 and 11. Trained model training and testing accuracies are shown for reference.
@@
We find that indeed accuracies of the MASC classifier on the random initialization outperforms the network, except for low corruption degrees (i.e. $<=20$% corruption degree). However, in the experiments where subspaces are trained on corrupted training data from corrupted models, by-and-large, the MASC classifier usually, and on at least one layer outperforms the MASC classifier trained on the random initialization with exceptions being the 80% corruption degree models on MLP-MNIST, AlexNet-Tiny ImageNet and 100% corruption degree on CNN-FashionMNIST.
@@
MASC accuracy on testing for random initialized model and trained model when data is projected on subspaces corresponding to true training labels is shown in Figure 10and 12. Notably, for the experiments where subspaces are constructed with true labels on corrupted models, the MASC classifier on these models outperforms the MASC classifier on random initializations usually and certainly in at least one layer on every model tested. These results are consistent with the main message of the paper, namely that even with memorized models, the layerwise representations of the models are organized in a manner that they develop significant ability to generalize over and above that bestowed by a random initialization, and in particular, they do not lose this ability, as one might have naively expected, due to label noise. If they were losing this ability, then the MASC classifier on the subspaces would end up performing significantly worse than the MASC classifier run on randomly initialized models.
@@
Although it is interesting that random projection have good generalization capabilities, it is not surprising as this has been shown by (Alain & Bengio, 2018) and studied by others (Jarrett et al., 2009).
@@
Figure 10: MASC accuracy over the layers of trained and random initialized network when the data set is projected onto subspace corresponding to true training labels. Testing accuracy of the trained model is shown for comparison.
@@
A.7 Minimum Angle Subspace Classifier Algorithm
For a given data point $\bm{x}$ from training or testing set, given a layer output data point $\bm{x_{l}}$ from layer $l$ when input $\bm{x}$ is passed through the network and its corresponding training subspaces $\{S_{k}\}_{k=1}^{K}$, we use Minimum Angle Subspace Classifier (MASC) Algorithm 1 for predicting class labels $y(\bm{x_{l}})$
For training dataset $\mathcal{D}\{(\bm{x_{i}},y_{i})\}_{i=1}^{m}\in\mathbb{R}^{d}\times\mathbb{R}$, where each $\bm{x_{i}}\in\mathbb{R}^{d}$ and $y_{i}\in\{C_{k}\}_{k=1}^{K}$ are input-label pairs, we estimate training subspaces $\{S_{k}\}_{k=1}^{K}$ for all classes $K$ and given layer $l$ of the neural network using Algorithm 2 and 3. For experiments in Section 4, Algorithm 2 uses corrupted training labels whereas in Section 5, Algorithm 2 uses true training labels to create class-specific subspaces.
Minimum Angle Subspace Classifier (MASC) Accuracy on Corrupted Training: After using Algorithm 1, the accuracy with respect to MASC predicted class labels of training data set and corrupted labels of training data are calculated.
Minimum Angle Subspace Classifier (MASC) Accuracy on Original Training: After using Algorithm 1, the accuracy with respect to MASC predicted class labels of training data set and true labels of training data is calculated.
Minimum Angle Subspace Classifier (MASC) Accuracy on Testing: After using Algorithm 1, the accuracy with respect to MASC predicted class labels of testing data set and original labels of testing data is calculated.
Algorithm 1 Minimum Angle Subspace Classifier (MASC)
1:Input: Training subspaces $\{S_{k}\}_{k=1}^{K}$, layer output data point $\bm{x_{l}}$ from layer $l$ when input $\bm{x}$ is passed through the network and classes $\{C_{k}\}_{k=1}^{K}$.
2:Output: MASC prediction class label $y(\bm{x_{l}})$ according to layer $l$
3:for each class $C_{k}$ do
4:     $\bm{x_{lk}}\longleftarrow\text{ compute the projection of }\bm{x_{l}}\text{ onto subspace }S_{k}$.
5:     Compute the angle $\theta(\bm{x_{l}},\bm{x_{lk}})$ between $\bm{x_{l}}$ and $\bm{x_{lk}}$
6:end for
7:Assign the label $y(\bm{x_{l}})=C_{k}$ where $k=\arg\min_{k}\theta(\bm{x_{l}},\bm{x_{lk}})$
Algorithm 2 Subspaces Estimator for MASC
1:Input: Training dataset $\mathcal{D}\{(\bm{x_{i}},y_{i})\}_{i=1}^{m}\in\mathbb{R}^{d}\times\mathbb{R}$, where each $\bm{x_{i}}\in\mathbb{R}^{d}$ and $y_{i}\in\{C_{k}\}_{k=1}^{K}$ are input-label pairs, neural network, and layer $l$.
2:Output: Subspaces $\{S_{k}\}_{k=1}^{K}$ for classes $K$ and given layer $l$.
3:$\mathcal{D}_{l}=\phi$
4:for each input pair $(\bm{x_{i}},y_{i})$ in $\mathcal{D}$ do Pass $\bm{x_{i}}$ through the network layers to obtain the output of layer $l$, denoted as $\bm{x_{l}}\in\mathbb{R}^{ld}$. $\mathcal{D}_{l}$= $\mathcal{D}_{l}\cup\{\bm{x_{l}}\}$
5:end for
7:Return: Subspaces $\{S_{k}\}_{k=1}^{K}$
Algorithm 3 PCA-Based Subspace Estimation
1:Input: Layer output $\mathcal{D}_{l}=\{(\bm{x_{l}},y_{i})\}_{i=1}^{m}$, where $\bm{x_{l}}\in\mathbb{R}^{ld}$ and $y_{i}\in\{C_{k}\}_{k=1}^{K}$
2:Output: Subspaces $\{S_{k}\}_{k=1}^{K}$ for classes $K$
3:$\mathcal{D}_{\text{new}}\leftarrow\mathcal{D}_{l}$
4:for each data point $\bm{x_{l}}\text{ in }\mathcal{D}_{l}$ do $\mathcal{D}_{\text{new}}\leftarrow\mathcal{D}_{\text{new}}\cup\{\bm{-x_{l}}\}$
5:end for
6:for each class $C_{k}$ in $C_{K}$ do
7:     Extract the subset of data $\mathcal{D}_{\text{new},k}=\{\bm{x_{l}}\mid y_{i}=k\}$
8:     Apply PCA to $\mathcal{D}_{\text{new},k}$ to calculate the PCA components
9:     The span of the PCA components defines the subspace $S_{k}$
10:end for
11:Return: Subspaces $\{S_{k}\}_{k=1}^{K}$
A.8 Experimental results on two additional models (MLP-CIFAR10 and AlexNet-Tiny ImageNet.)
All the experimental results on two additional models i.e, MLP-CIFAR10 and AlexNet-Tiny ImageNet are shown in this section.
MASC accuracy over the layers of the MLP trained on CIFAR10 and AlexNet trained on Tiny ImageNet when the data is projected onto corrupted training subspaces is shown in Figure 13. MASC accuracy over the layers of the MLP trained on CIFAR10 and AlexNet trained on Tiny ImageNet when the data set is projected subspace corresponding to true training labels is shown in Figure 14. MASC accuracy over the layers of the generalized MLP network trained on CIFAR10 and AlexNet network trained on Tiny ImageNet when the data is projected onto corrupted training subspaces is shown in Figure 15.
Figure 13: MASC accuracy over the layers of the network when the data is projected onto corrupted training subspaces with the indicated corruption degree. The number of class-wise PCA components of these models are shown in Figure 16 in section A.9 of the Appendix.
Figure 14: MASC accuracy over the layers of the network when the data set is projected subspace corresponding to true training labels. The respective number of class-wise PCA components for true training label subspaces of the models is shown in Figure 17 in section A.9 of the Appendix.
@@ A.7 Minimum Angle Subspace Classifier Algorithm
Figure 15: MASC accuracy over the layers of the generalized network when the data set is projected onto corrupted training subspaces with the indicated corruption degree. The respective number of class-wise PCA components of the models is shown in Figure 18 in Section A.9 of the Appendix.
A.9 Number of PCA components
@@
This section covers the number of class-wise PCA components used in all the experiments in the main paper.
@@
Number of class-wise PCA components of corrupted training subspace over the layer of MLP trained on MNIST and CIFAR-10, CNN trained on Fashion-MNIST, and AlexNet trained on CIFAR-100 and Tiny ImageNet is shown in Figure 16. For generalized models, it is observed that for 99% variance captured, the number of PCA components is significantly smaller in comparison to the ambient dimensionality of the layer (number of units in that layer). Over corruption, it is observed that for MLP-MNIST, MLP-CIFAR-10, and CNN-Fashion-MNIST, the number of class-wise PCA components increase. And the variance between the number of dimensions decrease. For AlexNet-CIFAR-100 and AlexNet-Tiny ImageNet, it is the opposite case, wherein the number of PCA components over corruption decreases.
@@
Figure 16: Class-wise number of PCA components of the corrupted training subspace used in section 4 over the layers of multiple networks with various corruptions degrees. Although it is not mentioned in the legend, all the 100 classes and 200 classes of CIFAR-100 and Tiny ImageNet respectively are plotted.
Number of class-wise PCA components of original training subspaces over the layer of networks used in Section 5 is shown in Figure 17. We find that for original training subspaces, although the dimensionality has increased with corruption degree, the variance has remained the approximately similar.
Figure 17: Class-wise number of PCA components of the subspace corresponding to true training labels used in section 5 over the layers of multiple networks with various corruptions. Although it is not mentioned in the legend, all the 100 classes and 200 classes of CIFAR-100 and Tiny ImageNet respectively are plotted.
Number of class-wise PCA components of original training subspaces over the layer of networks of the generalized model used in Section 6 is shown in Figure 18.
@@
Figure 18: Class-wise number of PCA components of the corrupted training subspace used in section 6 over the layers of multiple generalized networks with various corruption degrees. Although it is not mentioned in the legend, all the 100 classes and 200 classes of CIFAR-100 and Tiny ImageNet respectively are plotted.
A.10 SGD vs Adam
@@
For 40% corruption degree, although approximately 36% of labels are flipped, with MLP trained on MNIST, model trained on $Adam$ has MASC testing accuracy of around 95%, and around 50% for CIFAR-10. This is better than the MASC accuracy at the input layer and models testing accuracy. For 60% corruption degree with MLP trained on MNIST, model trained on $Adam$ has MASC testing accuracy of around 90% whereas for model trained on $SGD$ is around 85%. Although the networks are trained with 56% of label corruption, yet in FC (512) layer, the MASC training original accuracy is about 70% for model trained on $Adam$ whereas it is 85% for model trained on $SGD$. MASC on original training data does unfavorably on model trained with $SGD$ rather than with $Adam$, although further investigation is required.
Figure 19: MASC accuracy over the layers of the MLP network when the data is projected onto corrupted training subspaces with the indicated corruption degree, for MLP models with MNIST and CIFAR10 datasets. Rows corresponds to plots which have the same corruption degree and the columns correspond to the models with $SGD$ and $Adam$ optimizer as noted. Training and testing accuracy of the model is shown.FC corresponds to fully connected layer with ReLU activation.
Figure 20: Class-wise number of PCA components of the corrupted training subspace over the layers of MLP networks trained with MNIST and CIFAR10 datasets with various corruption degree. Rows corresponds to plots which have the same corruption degree and the columns correspond to the models with $SGD$ and $Adam$ optimizer as noted.
@@
Figure 21: MASC accuracy over the layers of the MLP network when the data set is projected onto subspace corresponding to true training labels. Rows corresponds to plots which have the same corruption degree and the columns correspond to the models with $SGD$ and $Adam$ optimizer as noted. Training and testing accuracy of the model is shown. FC corresponds to fully connected layer with $ReLU$ activation.
@@
Figure 22: Class-wise number of PCA components of the subspace corresponding to true training labels over the layers of MLP networks with various corruption degrees. Rows corresponds to plots which have the same corruption degree and the columns correspond to the models with $SGD$ and $Adam$ optimizer as noted.
@@
Figure 23: MASC accuracy over the layers of the generalized MLP network when the data set is projected onto corrupted training subspaces with the indicated corruption degree. Rows corresponds to plots which have the same corruption degree & the columns correspond to the generalized models with $SGD$ and $Adam$ as noted. Training & testing accuracy of the generalized model with $SGD$ and $Adam$ is shown. FC corresponds to fully connected layer with $ReLU$ activation.

### v2

@@ Appendix E Computational costs of MASC
Appendix F Experiments with AlexNet model trained on Tiny ImageNet
@@
MASC test accuracy over the layers of AlexNet trained on Tiny ImageNet when the data is projected onto corrupted training subspaces with the indicated corruption degree is shown in Figure 20. Test accuracy of the model and best model test accuracy is shown for comparison. Best model test accuracy corresponds accuracy of the test data of the model if early stopping was used.
@@
Even for AlexNet-Tiny ImageNet corrupted model (with non-zero corruption degree), except those with 100% corruption degree, we find that our Minimum Angle Subspace Classifier (MASC) in at least one layer has better test accuracy than the corresponding model itself.
@@
MASC test accuracy over the layers of AlexNet trained on Tiny ImageNet when the data set is projected onto corrupted training and true training subspace is shown in Figure 21.
Figure 20: MASC test accuracy over the layers of AlexNet trained on Tiny ImageNet when the data is projected onto corrupted training subspaces with the indicated corruption degree. Test accuracy of the model and best model test accuracy is shown for comparison. Best model test accuracy corresponds accuracy of the test data of the model if early stopping was used.
Figure 21: MASC test accuracy over the layers of AlexNet trained on Tiny ImageNet when the data set is projected onto corrupted training and true training subspace. Test accuracy of the model and best model test accuracy is shown for comparison. Best model test accuracy corresponds accuracy of the test data of the model if early stopping was used.
Appendix G Random initializations of control models versus trained models
@@
This section covers a set of control experiments to show MASC performance on random initialized model and contrast this with the trained models presented in the main text. We have verified that, for every such control model, the model training and test accuracies for the randomly initialized models is at chance level, for the corresponding dataset in question.
@@
MASC accuracy on test for randomly initialized model and trained model when data is projected on corrupted training subspaces is shown in Figure 22 and 24. Trained model training and test accuracies are shown for reference.
@@
We find that indeed accuracies of the MASC classifier on the random initialization outperforms the network, except for low corruption degrees (i.e. <=20<=20% corruption degree). However, in the experiments where subspaces are trained on corrupted training data from corrupted models, by-and-large, the MASC classifier usually, and on at least one layer outperforms the MASC classifier trained on the random initialization with exceptions being the 80% corruption degree models on MLP-MNIST, AlexNet-Tiny ImageNet and 100% corruption degree on CNN-Fashion-MNIST.
@@
MASC accuracy on test for random initialized model and trained model when data is projected on subspaces corresponding to true training labels is shown in Figure 23 and 25. Notably, for the experiments where subspaces are constructed with true labels on corrupted models, the MASC classifier on these models outperforms the MASC classifier on random initializations usually and certainly in at least one layer on every model tested. These results are consistent with the main message of the paper, namely that even with memorized models, the layerwise representations of the models are organized in a manner that they develop significant ability to generalize over and above that bestowed by a random initialization, and in particular, they do not lose this ability, as one might have naively expected, due to label noise. If they were losing this ability, then the MASC classifier on the subspaces would end up performing significantly worse than the MASC classifier run on randomly initialized models.
@@
Although it is interesting that random projection have good generalization to true labels, it is not surprising as this has been shown (Alain and Bengio, 2018) and also studied (Jarrett et al., 2009).
@@
Figure 23: MASC accuracy over the layers of trained and random initialized network when the data set is projected onto subspace corresponding to true training labels. Test accuracy of the trained model is shown for comparison.
@@
Appendix H Investigating latent memorization capabilities of uncorrupted models
@@
Figure 26: Minimum Angle Subspace Classifier (MASC) accuracy over the layers of the generalized network when the data set is projected onto corrupted training subspaces with the indicated corruption degree. Rows corresponds to plots which have the same corruption degree & the columns correspond to the generalized models as noted. Training & test accuracy of the generalized model is shown. FC corresponds to fully connected layer with ReLUReLU activation whereas Flat corresponds to flatten layer without ReLUReLU activation. The respective number of class-wise PCA components of the models is shown in Figure 34. SGDSGD optimizer was used for training MLP models, whereas AdamAdam optimizer was used for other models.
Conversely, we ask how well a network trained on true labels can manifest memorization of an arbitrary relabeling of its training data. More specifically, we built a MASC classifier on a model trained on true training labels, with the goal of memorizing training data whose labels are corrupted to varying degrees post hoc.
To do this, we shuffle the labels of the training set to some corruption degree and construct the corresponding class-specific subspaces with respect to the layerwise outputs of the generalized models, i.e. models trained with uncorrupted training data. We then build a MASC classifier corresponding to these subspaces.
MASC accuracy on corrupted training data, original training data and MASC accuracy on test data over the layers of the networks are shown in Figure 26. We have results on additional models, namely MLP trained on CIFAR-10, CNN trained on MNIST and CIFAR-10, and AlexNet trained on Tiny ImageNet; see Figure 29. In Table 5, we show by what percentage the MASC classifier outperformed the model train accuracies on the corrupted labels for the best layer for corruption degrees 20%, 40%, 60% and 80%.
Table 5: Percentage by which the MASC classifier (run on the best layer) outperformed the model’s train accuracy on corrupted labels. The model is trained on images with true training labels. MASC classifier is operating on subspaces corresponding to corrupted training data. The best layer corresponds to the one that has the highest measured MASC train accuracy among the layers for the said model/dataset. The accuracies in each case are averaged over three runs. The values are rounded to the second decimal place.
Corruption degree 20% 40% 60% 80%
MLP-MNIST -2.74% -2.35% -1.81% 3.11%
MLP-CIFAR-10 (Appendix) 18.4% 51.28% 112.81% 255.93%
CNN-MNIST (Appendix) -0.24% 0.40% 3.50% 26.76%
CNN-Fashion-MNIST -5.33% -1.01% 7.79% 50.05%
CNN-CIFAR-10 (Appendix) 13.17% 42.65% 101.38% 243.12%
AlexNet-CIFAR-100 27.91% 70.03% 152.78% 394.89%
AlexNet-Tiny ImageNet (Appendix) 23.19% 63.18% 144.52% 389.85%
Interestingly, we find a dichotomy in model behavior here, with some models trained on specific datasets having the propensity to memorize to a high degree, whereas others not demonstrating such ability. Specifically, we observe that for uncorrupted models with low/modest model test accuracies (i.e. AlexNet-CIFAR-100 and AlexNet-Tiny ImageNet), the MASC classifiers described above have high accuracies on the corrupted training set (i.e. appear to be able to memorize with high accuracies). Conversely, in most uncorrupted models with high model test accuracies (i.e. MLP-MNIST and CNN-Fashion-MNIST), we find that these MASC classifiers have more modest accuracies on the training set with corrupted labels (i.e. do not memorize to a high degree). This suggests the hypothesis that high generalization to true labels inhibits the innate ability of models to memorize and conversely that low/moderate generalization to true labels is accompanied by representations that offer greater propensity to memorize, at least as manifested by the MASC classifier. However, more work is needed to study this phenomenon to understand the trade-offs and the mechanisms underlying them, which has been beyond the scope of the present work. Also, surprisingly, we find that MASC classifiers often have test accuracies that approach or exceed uncorrupted model test accuracies, even though they correspond to corrupted subspaces (see e.g. AlexNet-CIFAR-100 and AlexNet-Tiny ImageNet).
Appendix I Experimental results on additional models (MLP-CIFAR-10, CNN-MNIST, CNN-CIFAR-10, AlexNet-Tiny ImageNet and ResNet-18-CIFAR-10)
In Table 6, we report the accuracy difference between the MASC classifier – when the data is projected onto corrupted training subspaces – and the model for the best such layer.
Table 6: Accuracy difference between the MASC classifier (run on the best layer) and the model’s test accuracy when the data is projected onto corrupted training subspaces. The model’s test accuracy are reported in Table 4. The best layer corresponds to the one that has the highest measured MASC test accuracy among the layers for the said model/dataset. The accuracies in each case are averaged over three runs. The values are rounded to the second decimal place. The values are rounded to the second decimal place.
Corruption degree 20% 40% 60% 80%
MLP-MNIST 9.55 23.18 40.65 51.33
MLP-CIFAR-10 (Appendix) 4.81 9.85 14.35 12.75
CNN-MNIST (Appendix) 8.59 25.72 46.49 56.91
CNN-Fashion-MNIST 5.82 20.55 36.74 48.33
CNN-CIFAR-10 (Appendix) 1.39 2.89 8.37 11.02
AlexNet-CIFAR-100 9.55 10.90 6.26 1.54
AlexNet-Tiny ImageNet (Appendix) 2.68 2.91 0.92 0.10
ResNet-18-CIFAR-10 -0.28 10.48 23.20 23.86
In Table 7, we show by what percentage the MASC classifier (subspace corresponding to true labels) outperformed the model for the best layer for corruption degrees 20%, 40%, 60% and 80%. In fact, the MASC test accuracies for the corrupted models (with non-zero corruption degree) are sometimes fairly close to the test accuracy of the uncorrupted model. In Table 8, we also report the accuracy difference between the MASC classifier – when the data is projected onto training subspaces corresponding to true labels – and the model for the best such layer.
Table 7: Percentage by which the MASC classifier (run on the best layer) outperformed the model’s test accuracy when the data is projected onto training subspaces corresponding to true labels. The best layer corresponds to the one that has the highest measured MASC test accuracy among the layers for the said model/dataset. The accuracies in each case are averaged over three runs. The values are rounded to the second decimal place.
Corruption degree 20% 40% 60% 80%
MLP-MNIST 10.96% 32.01% 77.77% 198.43%
MLP-CIFAR-10 (Appendix) 11.00% 30.27% 66.92% 146.09%
CNN-MNIST (Appendix) 10.69% 39.50% 104.97% 239.44%
CNN-Fashion-MNIST 8.37% 33.09% 88.97% 212.42%
CNN-CIFAR-10 (Appendix) 3.29% 13.43% 58.23% 138.65%
AlexNet-CIFAR-100 38.78% 68.91% 133.76% 337.51%
AlexNet-Tiny ImageNet (Appendix) 33.31% 83.85% 157.10% 212.33%
ResNet-18-CIFAR-10 2.92% 32.23% 93.55% 228.64%
Table 8: Accuracy difference between the MASC classifier (run on the best layer) and the model’s test accuracy when the data is projected onto training subspaces corresponding to true labels. The model’s test accuracy are reported in Table 4. The best layer corresponds to the one that has the highest measured MASC test accuracy among the layers for the said model/dataset. The accuracies in each case are averaged over three runs. The values are rounded to the second decimal place.
Corruption degree 20% 40% 60% 80%
MLP-MNIST 9.57 23.46 42.13 63.69
MLP-CIFAR-10 (Appendix) 5.35 12.21 20.44 28.76
CNN-MNIST (Appendix) 9.36 27.43 49.45 67.77
CNN-Fashion-MNIST 6.51 20.30 38.49 54.32
CNN-CIFAR-10 (Appendix) 1.99 6.20 18.03 25.40
AlexNet-CIFAR-100 11.03 14.15 12.90 11.58
AlexNet-Tiny ImageNet (Appendix) 3.25 4.57 3.17 1.55
ResNet-18-CIFAR-10 1.97 16.61 32.26 45.77
All the experimental results on additional models i.e, MLP-CIFAR-10, CNN-MNIST, CNN-CIFAR-10 and AlexNet-Tiny ImageNet are shown in this section.
MASC accuracy over the layers of the MLP trained on CIFAR-10, CNN trained on MNIST, CNN trained on CIFAR-10, and AlexNet trained on Tiny ImageNet when the data is projected onto corrupted training subspaces is shown in Figure 27. MASC accuracy over the layers of the MLP trained on CIFAR-10, CNN trained on MNIST, CNN trained on CIFAR-10, and AlexNet trained on Tiny ImageNet when the data set is projected subspace corresponding to true training labels is shown in Figure 28. MASC accuracy over the layers of the generalized MLP network trained on CIFAR-10, CNN trained on MNIST, CNN trained on CIFAR-10 and AlexNet network trained on Tiny ImageNet when the data is projected onto corrupted training subspaces is shown in Figure 29.
Figure 27: MASC accuracy over the layers of the network when the data is projected onto corrupted training subspaces with the indicated corruption degree. The number of class-wise PCA components of these models are shown in Figure 31.
Figure 28: MASC accuracy over the layers of the network when the data set is projected onto corrupted subspace and subspace corresponding to true training labels. The respective number of class-wise PCA components for true training label subspaces of the models is shown in Figure 33.
Figure 29: MASC accuracy over the layers of the generalized network when the data set is projected onto corrupted training subspaces with the indicated corruption degree. The respective number of class-wise PCA components of the models is shown in Figure 35.
Appendix J Number of PCA components
@@
This section covers the number of class-wise PCA components used in all the experiments.
@@
Number of class-wise PCA components of corrupted training subspace over the layer of MLP-MNIST,CNN-Fashion-MNIST, and AlexNet-CIFAR-100 is shown in Figure 30 and MLP-CIFAR-10,CNN-MNIST, CNN-CIFAR-10 and AlexNet-Tiny ImageNet is shown in Figure 31. For generalized models, it is observed that for 99% variance captured, the number of PCA components is significantly smaller in comparison to the ambient dimensionality of the layer (number of units in that layer). Over corruption, it is observed that for MLP-MNIST, MLP-CIFAR-10, and CNN-Fashion-MNIST, the number of class-wise PCA components increase. And the variance between the number of dimensions decrease. For AlexNet-CIFAR-100 and AlexNet-Tiny ImageNet, it is the opposite case, wherein the number of PCA components over corruption decreases.
@@
Figure 30: Class-wise number of PCA components of the corrupted training subspace over the layers of multiple networks with various corruptions degrees. Although it is not mentioned in the legend, all the 100 classes of CIFAR-100 are plotted.
@@
Figure 31: Class-wise number of PCA components of the corrupted training subspace over the layers of multiple networks with various corruptions degrees. Although it is not mentioned in the legend, all the 200 classes of Tiny ImageNet are plotted.
Number of class-wise PCA components of original training subspaces over the layer of networks is shown in Figure 32 and Figure 33. We find that for original training subspaces, although the dimensionality has increased with corruption degree, the variance has remained the approximately similar.
Figure 32: Class-wise number of PCA components of the subspace corresponding to true training labels over the layers of multiple networks with various corruptions. Although it is not mentioned in the legend, all the 100 classes of CIFAR-100 are plotted.
Figure 33: Class-wise number of PCA components of the subspace corresponding to true training labels over the layers of multiple networks with various corruptions. Although it is not mentioned in the legend, all the 200 classes of Tiny ImageNet are plotted.
Number of class-wise PCA components of original training subspaces over the layer of networks of the generalized model is shown in Figure 34 and Figure 35 .
Figure 34: Class-wise number of PCA components of the corrupted training subspace over the layers of multiple generalized networks with various corruption degrees. Although it is not mentioned in the legend, all the 100 classes of CIFAR-100 are plotted.
Figure 35: Class-wise number of PCA components of the corrupted training subspace over the layers of multiple generalized networks with various corruption degrees. Although it is not mentioned in the legend, all the 200 classes of Tiny ImageNet are plotted.
Appendix K SGD vs Adam
@@
For 40% corruption degree, although approximately 36% of labels are flipped, with MLP trained on MNIST, model trained on $Adam$ has MASC test accuracy of around 95%, and around 50% for CIFAR-10. This is better than the MASC accuracy at the input layer and models test accuracy. For 60% corruption degree with MLP trained on MNIST, model trained on $Adam$ has MASC test accuracy of around 90% whereas for model trained on $SGD$ is around 85%. Although the networks are trained with 56% of label corruption, yet in FC (512) layer, the MASC training original accuracy is about 70% for model trained on $Adam$ whereas it is 85% for model trained on $SGD$. MASC on original training data does unfavorably on model trained with $SGD$ rather than with $Adam$, although further investigation is required.
Figure 36: MASC accuracy over the layers of the MLP network when the data is projected onto corrupted training subspaces with the indicated corruption degree, for MLP models with MNIST and CIFAR-10 datasets. Rows corresponds to plots which have the same corruption degree and the columns correspond to the models with SGDSGD and AdamAdam optimizer as noted. Training and test accuracy of the model is shown.
Figure 37: Class-wise number of PCA components of the corrupted training subspace over the layers of MLP networks trained with MNIST and CIFAR-10 datasets with various corruption degree. Rows corresponds to plots which have the same corruption degree and the columns correspond to the models with SGDSGD and AdamAdam optimizer as noted.
@@
Figure 38: MASC accuracy over the layers of the MLP network when the data set is projected onto subspace corresponding to true training labels. Rows corresponds to plots which have the same corruption degree and the columns correspond to the models with SGDSGD and AdamAdam optimizer as noted. Training and test accuracy of the model is shown. FC corresponds to fully connected layer with ReLUReLU activation.
@@
Figure 39: Class-wise number of PCA components of the subspace corresponding to true training labels over the layers of MLP networks with various corruption degrees. Rows corresponds to plots which have the same corruption degree and the columns correspond to the models with SGDSGD and AdamAdam optimizer as noted.
@@
Figure 40: MASC accuracy over the layers of the generalized MLP network when the data set is projected onto corrupted training subspaces with the indicated corruption degree. Rows corresponds to plots which have the same corruption degree & the columns correspond to the generalized models with SGDSGD and AdamAdam as noted. Training & test accuracy of the generalized model with SGDSGD and AdamAdam is shown.